Grade 12Differential Calculus

show that the curves 4x =y^2 and 4xy=k cut at right angles,if k^2=512

taniska
11 Years agoGrade 12

1 Answer

Jitender Singh
11 Years ago
Ans:
Hello student,
Please find the answer to your question below


y^2 = 4x
Slope of tangent:
2yy' = 4
y' = \frac{2}{y}..............(1)
4xy = k
xy' + y = 0
y' = \frac{-y}{x}...................(2)
(1)\times (2) = -1
\Rightarrow x = 2
Put in the equation, we have
y = \sqrt{8}
k = 8\sqrt{8}
k^2 = 512